# Rethinking Prompt Optimizers: From Prompt Merits to Optimization

### v3

@@ Abstract
Zixiao Zhu1  Hanzhang Zhou1  Zijian Feng1  Tianjiao Li1 Chua Jia Jim Deryl2Mak Lee Onn2Gee Wah Ng2Kezhi Mao1thanks: Corresponding author 1School of Electrical and Electronic Engineering, Nanyang Technological University, Singapore 2Home Team Science and Technology Agency, Singapore {zixiao.zhu, hanzhang.zhou, zijian.feng, tianjiao.li, ekzmao}@ntu.edu.sg {Deryl_CHUA, MAK_Lee_Onn, NG_Gee_Wah}@htx.gov.sg
@@ 2.1 Merit Discovery
Prior studies explore various merits that contribute to prompt effectiveness. Wei et al. (2022) propose that including Chain-of-Thought (CoT) reasoning in prompts enhances performance, while Lampinen et al. (2022) find that Detailed Explanations improve prompt quality. Bsharat et al. (2023) introduce 26 prompting principles, categorized into dimensions such as Prompt Structure and Clarity, and Specificity and Information. Ye et al. (2024) propose three key components for complex reasoning prompts: Detailed Descriptions, Context Specification, and Step-by-Step Reasoning.
@@
Fig. 2 shows the top 10 merits most frequently associated with high-scoring prompts. Among them, several merits exhibit notably higher frequencies, revealing that high-quality prompts consistently reflect a set of core merits that can be effectively generated by lightweight LLMs. Motivated by this finding and prior insights, we propose the following design merits to guide the construction of optimal prompts:
@@ A B A Win Tie B Win A Win Tie B Win A Win Tie B Win $\Delta$ WR
MePO BPO 51.2 23.3 25.5 51.2 6.3 42.5 59.0 6.0 35.0 +19.5
MePO FIPO 45.2 32.3 22.5 49.0 11.8 39.2 49.9 11.4 38.7 +14.6
@@ 4.4 Ablation Study: Training Ablation
Tulu2-7b
-dpo
Training Free 49.58 34.30 32.93 40.96 72.60 46.22
@@
LLaMA2-
7b-chat-hf
Training Free 36.91 30.55 25.23 39.88 53.10 37.13
@@ 5 Related Work
Automatic Prompt Optimization (APO) provides a practical alternative to fine-tuning by optimizing prompts (Yang et al., ; Zhu et al., 2023). Beyond exploration-based methods search over predefined prompt pools (Ma et al., 2023; Shi et al., ; Zhu et al., 2025, 2024), recent works favor discrete exploitation-based methods that directly generate or revise prompts without relying on the quality of initial pools, making them more flexible and better suited for generative tasks beyond fixed-answer classification.
@@ 7 Limitations
Acknowledgements
The research is supported by the Home Team Science and Technology Agency.
@@ Appendix B Identifying Effective Prompt Merits via Comparative Evaluation
Prompt Effectiveness Evaluation.
@@
Prompt Preference Judgement.
Response Scoring.
The quality of the model’s response to each prompt is evaluated using DeepSeek-R1. The scoring template is displayed in Fig. 13, inspired by Yang et al. (2025).
Prompt Optimization.
@@ Appendix C POP Dataset Construction, Evaluation and Training Details
To simulate noisy user inputs—such as “how can i go fr sigapor?”—we degrade 10% of the raw prompts from the Alpaca and BPO datasets using a base model. The prompt used for degradation is shown in Fig. 7. Below is an example from the Alpaca dataset:
@@ E.4.1 Qualitative Analysis
E.4.2 Merit-Level Evaluation
@@ E.9 Case Study: Iterative Prompt Refinement
Who is considered the ’father’ of natural language processing (NLP)?
@@
Who is considered the ’father’ of natural language processing (NLP)?

### v4

@@ Abstract
Zixiao Zhu1   Hanzhang Zhou1   Zijian Feng1   Tianjiao Li1 Chua Jia Jim Deryl2Mak Lee Onn2Gee Wah Ng2Kezhi Mao1thanks: Corresponding author 1School of Electrical and Electronic Engineering, Nanyang Technological University, Singapore 2Home Team Science and Technology Agency, Singapore {zixiao001, hanzhang001, feng0119}@e.ntu.edu.sg {tianjiao.li, ekzmao}@ntu.edu.sg {Deryl_CHUA, MAK_Lee_Onn, NG_Gee_Wah}@htx.gov.sg
@@ 2.1 Merit Discovery
Literature Grounding: Prior studies explore various merits that contribute to prompt effectiveness. Wei et al. (2022) propose that including Chain-of-Thought (CoT) reasoning in prompts enhances performance, while Lampinen et al. (2022) find that Detailed Explanations improve prompt quality. Bsharat et al. (2023) introduce 26 prompting principles, categorized into dimensions such as Prompt Structure and Clarity, and Specificity and Information. Ye et al. (2024) propose three key components for reasoning prompts: Detailed Descriptions, Context Specification, and Step-by-Step Reasoning.
@@
As shown in Fig. 2, the merits most frequently associated with high-performing prompts include Clarity, Precision, and Conciseness. These results reveal that strong prompts consistently exhibit a core set of qualities that can be effectively generated even by lightweight LLMs. We further verify this trend using GPT-4o and Qwen3-32B (see Sec. B.2). Although Chain-of-Thought is frequently emphasized in the literature, it was less prominent in our dataset, likely due to its focus on general open-ended queries rather than reasoning-specific tasks. To strike a balance, we consolidate Conciseness and Chain-of-Thought into Concise Chain-of-Thought, mitigating verbosity while preserving reasoning cues and introduce Preserve Original Information to keep that essential content.
Motivated by empirical findings and prior insights, we propose the following design merits to guide the construction of optimal prompts:
@@ A B A Win Tie B Win A Win Tie B Win A Win Tie B Win $\Delta$ WR
MePO BPO 51.2 23.3 25.5 51.5 6.0 42.5 59.0 6.0 35.0 +19.5
MePO FIPO 45.2 32.3 22.5 49.0 11.5 39.5 49.9 11.4 38.7 +14.6
@@ 4.4 Ablation Study: Training Ablation
Tulu2-7b -dpo Training Free 49.58 34.30 32.93 40.96 72.60 46.22
@@
LLaMA2- 7b-chat-hf Training Free 36.91 30.55 25.23 39.88 53.10 37.13
@@ 5 Related Work
Automatic Prompt Optimization (APO) provides a practical alternative to fine-tuning by optimizing prompts (Yang et al., ; Zhu et al., 2023). Beyond exploration-based methods search over predefined prompt pools (Ma et al., 2023; Shi et al., ), recent works favor discrete exploitation-based methods that directly generate or revise prompts without relying on the quality of initial pools, making them more flexible and better suited for generative tasks beyond fixed-answer classification.
@@ 7 Limitations
Acknowledgments
This research was supported by the Home Team Science and Technology Agency (HTX), Singapore under the NTU-HTX collaboration project: Parsimonious Domain Specific Large Language Model Enabled Multimodality Sensemaking. We express our sincere appreciation to HTX for their continued support and collaboration.
@@ Appendix B Identifying Effective Prompt Merits via Comparative Evaluation
Prompt Preference Judgement.
Response Scoring.
The quality of the model’s response to each prompt is evaluated using DeepSeek-R1. The scoring prompt is shown in Fig. 15, inspired by Yang et al. (2025). This rubric ranges from 0 to 10, where a score difference of 4 or more points represents over half the scale and thus indicates a substantial quality gap. We consider responses scoring 5 or above as good, as they contain at least partially relevant information (e.g., “4–5: provides some relevant information but misses key aspects”). In contrast, responses rated 2–3 or lower often reflect minimal relevance or severe misconstrual of the ground truth. Accordingly, we define a “score difference greater than 4 points” as the threshold to identify prompts that elicit substantially better responses from those that do not.
Prompt Optimization.
B.1 Prompt Merits Recognition
@@
B.2 GPT-4o and Qwen3 Merits Recognition
To ensure that our results are not overly dependent on DeepSeek-R1, we additionally conducted merit discovery experiments with GPT-4o and Qwen3-32B [^9] https://huggingface.co/Qwen/Qwen3-32B on the same random sample of 2,000 prompt–response pairs, using the same discovery template (Fig. 10). The merits and their frequencies are given in Fig. 5 and Fig. 6, respectively.
Figure 5: Key merits of high-performing prompts extracted from GPT-4o evaluations.
Figure 6: Key merits of high-performing prompts extracted from Qwen3-32B evaluations.
Across DeepSeek-R1, GPT-4o, and Qwen3-32B, the top-3 merits (Clarity, Precision, Conciseness) consistently dominated. Although additional merits (e.g., Explicit Guidance, Conclusion) appeared in some models, they are less stable and not consistently aligned. This cross-model consistency strongly supports the robustness of our final merits identification.
@@ Appendix C POP Dataset Construction, Evaluation and Training Details
To simulate noisy user inputs—such as “how can i go fr sigapor?”—we degrade 10% of the raw prompts from the Alpaca and BPO datasets using a base model. The prompt used for degradation is shown in Fig. 9. Below is an example from the proposed dataset:
@@
Raw: Calculate the interior angle of a triangle when three of its sides are 8 cm, 3 cm, and 7 cm.
Degraded: find th angle of a traingle with sidz 8cm, 3cm, an 7cm.
Raw: Take this equation and solve it for x. 5x + 7 = 57
Degraded: solve 5x+7=57 what is x
Raw: What are the probabilities that I get struck by lightning
Degraded: What r the chances I get lightnin’?
@@ E.4.1 Qualitative Analysis
E.4.2 Real-World Cases
Raw Prompt: Is it safe to take Panadol with coffee?
MePO: Can you provide specific details on whether taking Panadol alongside coffee is considered safe for health? Please include any potential interactions or side effects to consider.
Raw Prompt: How can I reset a Singtel WiFi router safely?
MePO: What specific steps are required to reset a Singtel WiFi router, and what potential issues should be considered during the process?
Raw Prompt: What announcements did Apple make at the iPhone launch event this year?
MePO: Which key announcements did Apple release during its most recent iPhone launch event?
Raw Prompt: Do you think the man in the video is criminal? If so, what kind of crime did he commit? Explain your reasons step by step.
MePO: Evaluate whether the individual depicted in the provided video exhibits characteristics or engages in actions that suggest criminal behavior. If you determine that criminal activity is present, specify the type of crime and provide a detailed rationale for your assessment, including specific evidence from the video that supports your conclusion.
Raw Prompt: Did the first man provoke the second man, if so, how? Explain your reasons step by step.
MePO: Analyze whether the first individual intentionally provoked the second individual, providing specific details and reasoning for your conclusion step by step.
E.4.3 Merit-Level Evaluation
@@ E.5.1 Degraded Data Ablation
SOTA Comparison: We compared the MePO, BPO, and FIPO’s optimization ability under noisy raw prompts. Examples are shown below:
Raw Prompt: "python list sort unique value no dup how??"
MePO Prompt: How can you sort a Python list to ensure all values are unique and there are no duplicates?
FIPO Prompt: "How can I sort a list of unique values in Python, removing any duplicates?"
BPO Prompt: How can I sort a list of unique values in Python without any duplicates?
Raw Prompt: "Who the top NBA plyaers r in 2024"
MePO Prompt: Which specific NBA players are the top performers in the 2024 season?
FIPO Prompt: As of 2024, could you please provide a list of the top 10 NBA players based on their current performance, statistics, and overall impact on the game? This list should include players’ names, positions, and their respective teams. Additionally, could you also provide a brief analysis of what makes each player stand out in their respective roles? Please note that the ranking of these players is subject to change and may vary based on different sources and criteria. The information provided should be accurate as of 2024 and should not include any speculative or future projections beyond that year. Thank you for your assistance in providing this information."
BPO Prompt: Who are the top NBA players in 2024?
Raw Prompt: "idk whats the probablity of gettin hit by lightning lol"
MePO Prompt: What is the specific probability of being struck by lightning, and is this scenario considered likely or unlikely?
FIPO Prompt: "What is the average likelihood of a person being struck by lightning in their lifetime, and how can one assess this risk?"
BPO Prompt: What is the probability of being hit by lightning?
Raw Prompt: "last week u told me abt that machine learning thing, can u explain it again simpler"
MePO Prompt: Request for a simplified explanation of the machine learning topic you discussed with me last week.
FIPO Prompt: "Hey, could you please explain the concept of machine learning in simpler terms? Just to clarify, I’m asking about the process where a computer program learns from data and improves its performance over time, without being explicitly programmed. Could you also provide a basic example of how it works? Thank you!"
BPO Prompt: Can you explain machine learning and its applications in simpler terms?
Conclusion: These comparisons show that MePO produces concise, clarified prompts that remove noise while preserving intent. In contrast, BPO is closer to literal rewriting, and FIPO often produces verbose instructions.
@@ E.9 Case Study: Iterative Prompt Refinement
Who is considered the ‘father’ of natural language processing (NLP)?
@@
Who is considered the ‘father’ of natural language processing (NLP)?